Logiweb bibliography of check

Up Help

[0] check

[1] base

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-06-16.UTC:15:22:33.470061 = MJD-53902.TAI:15:23:06.470061 = LGT-4657188186470061e-6